OpenStudy (anonymous):

(sin(5x))^4=(3/8)-1/2cos(10x)+1/8cos(20x)

OpenStudy (anonymous):

I used half angle identity. \[\sin(a/2)=\pm \sqrt(\frac{ 1-\cos(a) }{ 2 }\]

OpenStudy (anonymous):

so my question is how would you solve (sin(5x))^4 to get that answer

OpenStudy (anonymous):

could take the sqrt(1-cos(a)/2)^4 and A would be 10x

OpenStudy (anonymous):

because were taking it to an even power (4) the +/- would just become + and the -cos(a) would become +cos(a)

OpenStudy (anonymous):

not solving for x just need help proving it.